# Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training

### v4

@@ 3.3 Considering the Model Size to Obtain the Size-dependent MI metric
The MI metric does not account for the impact of model size. Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger [17]. In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters [16]. Therefore, we propose to enhance the MI metric by incorporating model size $\Phi$, measured in billions of parameters.
@@
Based on these principles, we propose the SMI metric—an exponential function with the normalized MI as the base and $1+\frac{1}{\Phi}$ as the exponent.
$SMI(s,o,\Phi)=Norm(MI(s,o))^{1+\frac{1}{\Phi}}.$ (4)
In the above formula, the MI metric quantifies the information of knowledge triples in the pre-training data, while the model size $\Phi$ implies the memory capacity of LLMs. These two factors jointly determine the model’s knowledge retention. We normalize the MI metric from its original range of -30 to 0 onto a 0–1 scale. Since the normalized MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
Figure 3: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families. Model size is reported in billions of parameters and plotted on a base-2 logarithmic scale. Variations in pre-training data across model families may account for differences in performance.
Table 1: R² and MSE (×$\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters. Across all models, MI consistently outperforms CO-OCCUR, and SMI further improves upon MI.
Model R² ($\uparrow$ higher is better) MSE (×$\times 10^{-2}$, $\downarrow$ lower is better)
@@
bloom-560m 0.384 0.589 0.624 8.566 7.729 7.652
bloom-1b1 0.577 0.715 0.756 9.123 7.950 7.824
bloom-1b7 0.584 0.740 0.772 9.389 8.083 7.970
bloom-3b 0.705 0.782 0.807 9.314 7.863 7.764
bloom-7b1 0.676 0.804 0.814 9.943 8.309 8.263
bloom (176b) 0.644 0.764 0.764 7.327 6.029 6.029
@@
pythia-14m 0.240 0.375 0.496 6.855 7.183 6.978
pythia-31m 0.287 0.381 0.549 6.905 7.202 6.905
pythia-70m 0.355 0.430 0.585 6.749 7.051 6.783
pythia-160m 0.454 0.522 0.611 6.877 6.969 6.796
pythia-410m 0.524 0.701 0.759 7.439 7.035 6.889
pythia-1.4b 0.574 0.799 0.835 8.254 7.325 7.202
pythia-2.8b 0.631 0.845 0.866 8.593 7.505 7.421
pythia-6.9b 0.637 0.855 0.865 8.538 7.373 7.334
pythia-12b 0.602 0.853 0.858 8.741 7.514 7.494
@@
ours-1.6b 0.710 0.794 0.848 8.563 7.696 7.462
ours-7b 0.784 0.847 0.855 8.983 8.033 7.997
ours-13b 0.661 0.816 0.820 8.585 7.688 7.674
4 Experiments
In the Experimental Setup section, we begin by cleaning and enhancing an open-source Wikipedia knowledge triple dataset to make it suitable for inference with pre-trained models. We then introduce five large-scale open-source pre-training corpora and perform retrieval-based analyses on them. This is followed by a description of five families of pre-trained language models and the corresponding evaluation protocol. In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across three evaluation metrics. We also conduct a hyperparameter study to further explore model behavior under different configurations.
4.1 Experimental Setup
Evaluation data
We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours [9]. We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2. For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
Pre-training corpora
  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs. It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality [11].
  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project. It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency [18, 30].
  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs [32].
  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes. It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM [26].
  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency [10].
Pre-training data retrieval
We process the pre-training corpus at the document level. For each knowledge triple, we retrieve the frequency of occurrences of the subject, the object, and their co-occurrences. Using the retrieval results from all documents, we calculate the co-occurrence metric and MI metric for each knowledge triple. Subsequently, we compute the SMI metric based on the MI metric and the model size. The retrieval process runs on 100 CPUs for 3 days.
Pre-trained models
We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs [34]. These models not only make the model weights available but also provide complete pre-training data. They are primarily categorized into the following families:
  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion [30].
  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile [4, 5, 35].
  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters. These models use a uniform data and training process, facilitating research that is both replicable and interpretable [3].
  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata. It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions [41, 32, 20]. The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
Evaluation setting
We assess these pre-trained models using the designated evaluation dataset. During the evaluation process, we set the temperature to 0.7, the maximum new tokens to 32, and the precision to bfloat16. We consider a response to be correct if the object appears within it. To ensure stability in the evaluation, we perform inference 20 times for each of the 20 queries in the QA task, subsequently calculating the average accuracy of 400 responses. This inference process runs for 5 days on 8 H20 GPUs.
The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2). The majority of QA data co-occur between 10 and 1000 times. To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values. For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2. The SMI metric is computed based on the MI metric.
@@ 4.1 Experimental Setup
Figure 4: Relationship between the SMI metric and average accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
(a) $\text{R}^{2}$ on smaller models
(b) $\text{R}^{2}$ on larger models
Figure 5: Comparison of $\text{R}^{2}$ results for three metrics—CO-OCCUR, MI, and SMI—across models of different sizes. (a) Smaller models (bloom-560m, gpt-neo-1.3B, pythia-14m, ours-1.6b); (b) Larger models (bloom, gpt-neox-20b, pythia-12b, ours-13b). $\uparrow$ Higher is better.
Table 2: Hyperparameter analysis of the SMI metric across different BLOOM models. (a) illustrates the impact of varying unit scales, while (b) shows the effect of different pre-normalization ranges.
(a) SMI values across varying unit scales
Unit bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) $10^{8}$ 0.590 0.716 0.741 0.783 0.804 0.764 $10^{9}$ 0.600 0.723 0.745 0.785 0.805 0.764 $10^{10}$ 0.624 0.756 0.772 0.807 0.814 0.764 $10^{11}$ 0.394 0.497 0.586 0.760 0.826 0.766 $10^{12}$ 0.343 0.222 0.197 0.235 0.349 0.769
(b) SMI values under different pre-normalization ranges
Range bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) [30,0][-30,0][ - 30 , 0 ] 0.624 0.756 0.772 0.807 0.814 0.764 [35,0][-35,0][ - 35 , 0 ] 0.626 0.750 0.766 0.800 0.811 0.764 [40,0][-40,0][ - 40 , 0 ] 0.625 0.746 0.761 0.797 0.809 0.764 [45,0][-45,0][ - 45 , 0 ] 0.623 0.742 0.758 0.794 0.808 0.764 [50,0][-50,0][ - 50 , 0 ] 0.620 0.739 0.756 0.793 0.808 0.764
@@ 4.2 Experimental Results
Results of SMI, MI, and co-occurrence metrics
The primary experimental results for the three metrics are presented in Table 1. Overall, we observe the following: (1) The SMI metric performs exceptionally well on the evaluation set, with both R² and MSE surpassing the baseline across all models. (2) For all models, the SMI metric achieves R² values greater than 0.49, and for models with parameters exceeding 1 billion, the R² exceeds 0.75. This highlights a robust linear relationship and effectively predicts QA accuracy. (3) The MI metric outperforms the baseline co-occurrence metric, while the SMI metric further exceeds the MI metric, demonstrating the effectiveness of our iterative improvement approach.
The comparison of experimental results from models of different sizes indicates diminishing returns when increasing model size. A linear relationship is observed between the models’ average accuracy on the full evaluation dataset and the logarithm of model size, suggesting that significant gains in QA performance require exponential increases in model size (Figure 3). Meanwhile, the advantage of the SMI metric over the MI metric is more pronounced in smaller models (Figure 5). As model size increases, this advantage becomes less evident. For the BLOOM 560m model, the R² values differ by 3.55E-02, and the MSE values differ by 7.75E-04. In contrast, for the BLOOM 176b model, the difference in R² values is only 2.42E-04, and the difference in MSE values is just 6.60E-06. This indicates that larger models may possess sufficient memory capacity.
Moreover, the benefits of data optimization exhibit diminishing returns and an upper limit. Although a higher SMI generally correlates with better performance, achieving further gains becomes increasingly difficult at advanced levels due to the logarithmic growth of SMI (Figure 4). Specifically, the upper accuracy limits for the models bloom, gpt-neox-20b, pythia-12b, and ours-13b are 0.638, 0.727, 0.75, and 0.771, respectively. This indicates that even with perfect optimization of the pre-training data for these models, the accuracy on the QA task cannot exceed an approximate upper limit of 0.8.
Results of hyperparameter influence on the SMI metric
The SMI framework involves two critical hyperparameters. The first is the unit for model parameter quantity, which is set by default to 1e9. The second is the pre-normalization range for the SMI metric, with a default setting of -30 to 0. We conduct ablation experiments on these two hyperparameters, and the results are presented in Table 2. From the experimental results, we observe the following: (1) The unit for model parameter quantity significantly affects the SMI results. Using 1e10 as the unit yields the best overall performance. However, as the model size increases, the optimal unit value also increases. Smaller models exhibit lower robustness to parameter variations, whereas larger models display enhanced robustness. (2) The pre-normalization range has a relatively minor effect on the SMI results. The range of -30 to 0 performs best overall. Although a smaller pre-normalization range is preferable, the minimum value in the experimental data approximates -30, suggesting that further reduction might result in data loss.
5 Limitations and Future Work
Although the SMI metric shows strong predictive power for model accuracy on closed-book QA, the current evaluation is limited to this task type. This focus is intentional, as closed-book QA serves as a direct probe of factual knowledge internalized during pre-training, rather than general reasoning or transfer ability. However, broader task coverage is necessary to assess the general applicability of SMI across diverse task formats. Some variation is also observed across model scales, suggesting that future formulations should account for size-invariance. In addition, while SMI exhibits empirical effectiveness, its mathematical foundations remain underexplored. Future work will aim to establish a more rigorous theoretical grounding by linking information-theoretic principles with LLM behavior.
@@ 6 Conclusion
In this paper, we explore how to predict model knowledge using only pre-training signals and propose effective solutions to address three major challenges. We introduce SMI, an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy without any additional training. SMI achieves $R^{2}>0.75$ on models with over one billion parameters. Experimental results further reveal that both scaling model size and optimizing data yield diminishing returns, with QA accuracy plateauing at around 80% under optimal pre-training conditions. This indicates an upper bound on factual knowledge retention during pre-training, highlighting the need for downstream augmentation such as retrieval or prompting.
This paper emphasizes the feasibility of predicting downstream performance based on pre-training signals. It offers a principled metric for data-model-task alignment, thereby providing a reference for optimizing data selection and performance prediction during the pre-training phase of LLMs. To facilitate further research, all evaluation data and analysis code are made publicly available.
Appendix A Licenses and Terms of Use

### v5

@@ 3.3 Considering the Model Size to Obtain the Size-dependent MI metric
The MI metric does not account for the impact of model size. Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger (Kaplan et al., 2020). In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters (Kandpal et al., 2023). Therefore, we propose to extend the MI metric by explicitly incorporating the parameter count $\Phi$ as a measure of model size.
@@
Since linear transformations do not alter the strength of linear relationships, an exponential formulation offers a simple yet effective way to incorporate model size while capturing non-linear effects. Based on these principles, we propose the SMI metric, an exponential function with the MI metric as the base and $1+\frac{1}{\log\Phi}$ as the exponent.
$SMI(s,o,\Phi)=MI(s,o)^{1+\frac{1}{\log\Phi}}.$ (4)
In the above formula, the MI metric quantifies the information of knowledge triples in the pre-training data, while the model size $\Phi$ implies the memory capacity of LLMs. These two factors jointly determine the model’s knowledge retention. Since the MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
4 Experiments
In the Experimental Setup section, we begin by cleaning and enhancing an open-source Wikipedia knowledge triple dataset to make it suitable for inference with pre-trained models. We then introduce five large-scale open-source pre-training corpora and perform retrieval-based analyses on them. This is followed by a description of five families of pre-trained language models and the corresponding evaluation protocol. In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across metrics.
4.1 Experimental Setup
Evaluation data
We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours (Elazar et al., 2021). We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2. For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
Pre-training corpora
  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs. It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality (Gao et al., 2021).
  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project. It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency (Laurençon et al., 2022; Scao et al., 2022).
  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs (Soboleva et al., 2023).
  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes. It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM (Penedo et al., 2023).
  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency (Foundation, ).
Pre-training data retrieval
We process the pre-training corpus at the document level. For each knowledge triple, we retrieve the frequency of occurrences of the subject, the object, and their co-occurrences. Using the retrieval results from all documents, we calculate the co-occurrence metric and MI metric for each knowledge triple. Subsequently, we compute the SMI metric based on the MI metric and the model size. The retrieval process runs on 20 CPUs for 15 days.
Table 1: R² and MSE ($\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters. Across all models, MI consistently outperforms CO-OCCUR, and SMI further improves upon MI.
Model R² ($\uparrow$ higher is better) MSE ($\times 10^{-2}$, $\downarrow$ lower is better)
@@ 4.1 Experimental Setup
bloom-560m 0.384 0.589 0.591 8.566 7.729 7.723
bloom-1b1 0.577 0.715 0.719 9.123 7.950 7.939
bloom-1b7 0.584 0.740 0.743 9.389 8.083 8.072
bloom-3b 0.705 0.782 0.786 9.314 7.863 7.849
bloom-7b1 0.676 0.804 0.807 9.943 8.309 8.296
bloom 0.644 0.764 0.766 7.327 6.029 6.026
@@
pythia-14m 0.240 0.375 0.378 6.855 7.183 7.177
pythia-31m 0.287 0.381 0.384 6.905 7.202 7.196
pythia-70m 0.355 0.430 0.434 6.749 7.051 7.044
pythia-160m 0.454 0.522 0.526 6.877 6.969 6.961
pythia-410m 0.524 0.701 0.705 7.439 7.035 7.024
pythia-1.4b 0.574 0.799 0.803 8.254 7.325 7.312
pythia-2.8b 0.631 0.845 0.848 8.593 7.505 7.491
pythia-6.9b 0.637 0.855 0.859 8.538 7.373 7.360
pythia-12b 0.602 0.853 0.856 8.741 7.514 7.503
@@
ours-1.6b 0.710 0.778 0.782 8.563 7.712 7.694
ours-7b 0.784 0.837 0.839 8.982 8.028 8.016
ours-13b 0.661 0.807 0.810 8.584 7.674 7.665
@@
Figure 3: Relationship between the SMI metric and accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
Figure 4: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families.
Pre-trained models
We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs (Vaswani et al., 2017). These models not only make the model weights available but also provide complete pre-training data. They are primarily categorized into the following families:
  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion (Scao et al., 2022).
  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile (Black et al., 2021, 2022; Wang, 2021).
  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters. These models use a uniform data and training process, facilitating research that is both replicable and interpretable (Biderman et al., 2023b).
  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata. It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions (Zhang et al., 2024; Soboleva et al., 2023; Li et al., 2023). The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
Evaluation setting
We evaluate these pre-trained models on the designated QA evaluation dataset. During inference, we configure the generation parameters with a temperature of 0.7, a maximum of 32 new tokens, and bfloat16 precision. A model response is considered correct if the expected object entity appears anywhere within the generated text. To obtain stable and reliable estimates of performance, we repeat the inference procedure 20 times for each of the 20 queries in the QA task, resulting in a total of 400 responses per model. The final accuracy is computed as the mean over all 400 responses. The full evaluation runs for approximately five days on 8 H100 GPUs.
The retrieval results from the pre-training corpus show that the co-occurrence frequency distribution of question and answer entities is highly skewed (see Figure 2). Most QA pairs co-occur between 10 and 1000 times, while only a small portion appears at very low or very high frequencies. To mitigate this imbalance and enable meaningful evaluation across different frequency ranges, we group QA data with similar metric values. For the co-occurrence and MI metrics, we compute average accuracy by binning values into intervals of 0.2 on the unnormalized scale. The SMI metric is then derived from the MI metric after binning.
@@ 4.2 Experimental Results
Comparison of SMI, MI, and CO-OCCUR
Table 1 presents the R² and MSE results of CO-OCCUR, MI, and SMI across 24 models, with sizes ranging from 14M to 176B parameters. Overall, MI consistently improves over the co-occurrence baseline. For example, on gpt-neo-2.7B, R² increases from 0.672 (CO-OCCUR) to 0.829 (MI), while MSE decreases from 8.168 to 7.307. This suggests that jointly modeling frequency and specificity provides more informative signals for predicting closed-book QA performance. Incorporating model size into the metric (SMI) yields further gains across all models, though the magnitude of improvement is typically moderate. For instance, on pythia-12b, R² rises from 0.853 (MI) to 0.856 (SMI), and on bloom, from 0.764 to 0.766. These results suggest that incorporating model size into the metric provides a more principled account of the variance in observed QA accuracy.
Diminishing returns from data scaling
Despite the improved predictive power of SMI, we find that knowledge retention is limited by the nature of pre-training data. As shown in Figure 3, even when SMI reaches its maximum value of $1$, the closed-book QA accuracy does not exceed moderate levels. Specifically, the empirical upper bounds of accuracy for bloom, gpt-neox-20b, pythia-12b, and ours-13b are approximately $0.527$, $0.578$, $0.588$, and $0.630$, respectively. These results indicate that simply scaling the data cannot indefinitely improve factual recall.
Diminishing returns from model scaling
We further analyze the impact of model size on knowledge retention. Figure 4 shows that the average accuracy on the full evaluation set scales approximately linearly with the logarithm of model parameters, which is consistent with prior scaling-law observations. However, the growth rate becomes increasingly shallow for very large models: even at 176176B parameters, the accuracy remains below $0.2$. This finding suggests that model scaling alone is insufficient for achieving high factual coverage and further motivates the exploration of complementary approaches, such as retrieval-augmented generation or external memory modules.
@@ 5 Conclusion
In this work, we introduce a framework for quantifying knowledge retention in pre-trained language models, defined as the probability that a model reproduces factual information from its training corpus in a closed-book setting. We propose SMI, an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size, and demonstrate its effectiveness in forecasting closed-book QA accuracy. Our large-scale evaluation across 24 models shows that SMI achieves high predictive correlation (R² > 0.7 for models above 1B parameters) and remains robust across model families and domains.
Beyond prediction, our analysis reveals diminishing returns from scaling data and model size, suggesting an upper bound on factual knowledge that can be stored through pre-training alone. At the same time, SMI provides a diagnostic lens for data and model design: increasing the frequency of subject–object co-occurrences can strengthen knowledge memorization, filtering out irrelevant material can improve specificity, and scaling model size can further enhance retention, albeit with diminishing returns. These insights suggest that complementary strategies such as retrieval-augmented generation, external memory modules, and targeted fine-tuning may be necessary to further improve factual recall in LLMs.
We hope that our work provides a foundation for more systematic investigations into the interaction between corpus composition, model size, and factual knowledge retention, and informs future efforts to build models that combine parametric and non-parametric memory in a principled manner.
6 Limitations
While our study provides an information-theoretic framework for predicting knowledge retention, several limitations remain. First, our analysis is restricted to closed-book QA tasks and does not cover open-ended generation, reasoning-intensive tasks, or multi-hop factual queries, which may involve different memorization dynamics. Second, we rely on publicly available corpora and approximate document matching, which may introduce noise when measuring fact frequency and specificity. Third, our experiments focus on models with up to tens of billions of parameters, leaving open the question of whether the observed trends hold for frontier-scale models with hundreds of billions of parameters. Finally, our approach does not address how training dynamics, such as optimization schedules or curriculum learning, may interact with knowledge retention, which we leave for future work.
Appendix A AI Assistants in Research or Writing
In preparing this manuscript, AI assistants were employed solely to assist with refining the clarity, style, and readability of certain text segments. They were not involved in designing the study, developing or implementing the methodology, collecting or analyzing data, or generating the primary scientific contributions. All substantive research decisions, analyses, and conclusions are fully the responsibility of the authors.
Appendix B Licenses and Terms of Use